# Grounding Foundation Models through Federated Transfer Learning: A General Framework

### 10 Dec 2023 v5

@@ 3.2. Formulations of Machine Learning Tasks
where $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is the local FM of client $k$, $\phi_{\mathcal{W}^{G}_{a}}$ is the global PEFT module (e.g., prompt model, adaptor, and LoRA) parameterized by $\mathcal{W}^{G}_{a}$, and $\oplus$ denotes an operator that composes $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ with PEFT module $\phi_{\mathcal{W}^{G}_{a}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training the global PEFT module $\phi_{\mathcal{W}^{G}_{a}}$ based on client $k$’s local data while keeping $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ frozen.
The research works, FedPETuning (Zhang et al., 2023c), FedPrompt (Zhao et al., 2023b), FedAdapter (Cai et al., 2023b), and FedIT (Zhang et al., 2023b) adhere to the formulation of Eq.(13). FedPrompt, FedAdapter, and FedIT specifically investigate the impact of prompt tuning, adapter, and LoRA under federated learning, respectively. On the other hand, FedPETuning delves into the exploration of four PEFT techniques, namely adapter tuning, prefix tuning, LoRA, and BitFit, while operating under the FedAvg protocol.
@@ 4. FTL-FM Taxonomy
Semi-Honest $\mathcal{D}^{S}$ ($\mathcal{W}^{S}$) PPDG  (Carranza et al., 2023; Kurakin et al., $*{}^{*}$  (Mattern et al., 2022; Yue et al., $*{}^{*}$ DP  (Behnia et al., 2022; Bu et al., 2022; Shi et al., $*{}^{*}$, MP (Youssef et al., $*{}^{*}$ DP  (Yu et al., 2022; Mireshghallah et al., $*{}^{*}$ CP  (Xiao et al., 2023; Fan et al., ${}^{\dagger}$  (Chua et al., 2023; Khalid et al., ${}^{\dagger}$
$\mathcal{D}^{C}$ ($\mathcal{W}^{C}$) DP (Li et al., $*{}^{*}$, WPE (Xu et al., $*{}^{*}$, PTB (Zhou et al., $*{}^{*}$ PPDS (Hou et al., 2023b; Yu et al., $*{}^{*}$ DP (Wang et al., $*{}^{*}$ SA (WeBank, [n. $*{}^{*}$ DP (Azam et al., 2023; Balunovic et al., $*{}^{*}$, FE (Gupta et al., $*{}^{*}$ HE (Jin et al., $*{}^{*}$ DP  (Xu et al., 2023a; Zhao et al., $*{}^{*}$, PEFT (Zhang et al., $*{}^{*}$, SA (Fan et al., $*{}^{*}$
@@
5. Foundation Models for Enhancing Domain Models
In this section, we overview FTL-FM methods that fall into setting 1⃝ (see Definition 3.1), the objective of which is to transfer and adapt the knowledge of foundation models (FMs) to downstream clients’ domain models (DMs) for exploiting the power of FMs. To this end, the literature proposed a variety of methods involving data-level knowledge transfer, representation-level knowledge transfer, and model-level knowledge transfer. Table 3222While we consider an FTL-FM method should leverage certain protection methods to protect data or models, some highly relevant works focus on designing knowledge transfer methods without considering data or model protection. We still overview these methods and include them in the Table for completeness (same for Section 6 and Section 7). These methods leave ”blank” in the ”What To Protect” and ”How To Protect” columns. summarizes these methods.
@@ 5. Foundation Models for Enhancing Domain Models
BBT (Sun et al., 2022b, a) $\hat{Y}$ of $\mathcal{W}^{S}_{\text{FM}}$ $X$ + soft $P$
@@
BDPL (Diao et al., 2023) $\hat{Y}$ of $\mathcal{W}^{S}_{\text{FM}}$ $X$ + hard $P$
Fed-BBPT (Lin et al., 2023) Fr. $S_{\text{FM}}$: $\hat{Y}$ of $\mathcal{W}^{S}_{\text{FM}}$ Fr. $S_{\text{Fed}}$: Global model To $S_{\text{FM}}$: $X$ + soft $P$ To $S_{\text{Fed}}$: $\mathcal{W}^{C}$
FedBPT (Sun et al., 2023) Fr. $S_{\text{FM}}$: $\hat{Y}$ of $\mathcal{W}^{S}_{\text{FM}}$ Fr. $S_{\text{Fed}}$: Global distribution of $P$ To $S_{\text{FM}}$: $X$ + soft $P$ To $S_{\text{Fed}}$: Local distribution of $P$
ZooPFL (Lu et al., 2023) Fr. $S_{\text{FM}}$: $\hat{Y}$ of $\mathcal{W}^{S}_{\text{FM}}$ Fr. $S_{\text{Fed}}$: Global encoder To $S_{\text{FM}}$: Perturbed $X$ To $S_{\text{Fed}}$: Encoder
@@
FedBERT (Tian et al., 2022) Backward gradient Repr
@@
GPT-FL (Zhang et al., 2023a) Model trained on data generated by $\mathcal{W}^{S}_{\text{FM}}$
@@
BD-FMFL (Li et al., 2023d) BD global model
@@
5.1. FM adapted to DMs through data-level knowledge transfer
The literature has explored mainly two categories of methods that adapt the knowledge of FMs to specific domains through data-level knowledge transfer: (1) An FM server transfers data generated by general-purpose FMs or FMs augmented with industry-level knowledge (e.g., LawGPT (Nguyen, 2023), FinGPT (Yang et al., 2023), BloombergGPT (Wu et al., 2023a), and Med-PaLM2 (Singhal et al., 2023)) to clients of specific domains; (2) downstream clients prompt tuning a closed-source FM (e.g., ChatGPT) hosted by a server by training their local prompt generators based on domain-specific data and predictions generated by the FM. Optionally, these clients can enhance their local prompt generators through federated learning. The two categories of methods are illustrated in Figure 5(a) and Figure 5(b), respectively.
@@ 5.1. FM adapted to DMs through data-level knowledge transfer
Figure 5. Adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs) through data-level knowledge transfer.
Alternatively, clients can adapt knowledge of the FM hosted by an FM server to their specific domains through prompt tuning. For example, BBT (Sun et al., 2022b), BBTv2 (Sun et al., 2022a), and BlackVIP (Oh et al., 2023) proposed that a client can customize a remote FM (e.g., ${\text{RoBERTa}}_{\text{large}}$ (Liu et al., 2019), ${\text{BART}}_{\text{large}}$ (Lewis et al., 2020), and ViT-B (Radford et al., 2021)) by optimizing a domain-specific prompt generator based on its local data together with the predictions (e.g., logits) generated by the remote FM. BBT and BBTv2 employ a gradient-free method called CMA-ES (Covariance Matrix Adaptation Evolution Strategy) (Hansen, 2016) to optimize the client’s local prompt generator, whereas BlackVIP leverages SPSA (Simultaneous Perturbation Stochastic Approximation) (Spall, 1992) to efficiently approximates the high-dimensional gradients, which then are used to optimize the prompt generator. BBT, BBTv2, and BlackVIP require the FM API to take continuous (i.e., soft) prompts as input, and thus, they fail to be applied to FM APIs that only accept discrete inputs. To eliminate this constraint, BDPL (Diao et al., 2023) was proposed to leverage a variance-reduced policy gradient algorithm (Zhou et al., 2021) to optimize a discrete (i.e., hard) prompt generator at the client side. BBT, BBTv2, BlackVIP, and BDPL involve only one client, and thus, they correspond to stage 2 of the first objective formulated in Eq.(2) with $K=1,p^{C_{1}}=1$, $\eta=0$, the domain model $\mathcal{W}^{C_{1}}$ is the client’s local prompt generator and the specific form of the loss $\ell^{C_{1}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{1}};d)$ is formulated in Eq.(7).
Fed-BBPT (Lin et al., 2023) and FedBPT (Sun et al., 2023) extend the one-client setting to multiple clients that collaboratively train a global prompt generator through federated learning. More specifically, both Fed-BBPT and FedBPT involve an FM server $S_{\text{FM}}$ hosting an FM and a federated server $S_{\text{Fed}}$ for aggregating clients’ local prompt generators, as illustrated in Figure  5(b). In Fed-BBPT, each client leverages the Simultaneous Perturbation Stochastic Approximation (Spall, 1997) to train its prompt generators using local data and predictions generated by the FM, and the federated server forms the global model by aggregating clients’ prompt generators. FedBPT differs from Fed-BBPT in that each client in FedBPT exploits CMA-ES to optimize the distribution of prompts generated by its prompt generator, and the server derives the global prompt generator by aggregating local distributions of clients’ prompts. Generally, Fed-BBPT and FedBPT adhere to the objective formulated in Eq.(2) with $K>1$, $\eta=0$, the domain model $\mathcal{W}^{C_{k}}$ of each client $k$ is a prompt generator and the specific loss of federated prompt tuning $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ between the FM server and a client $k$ is formulated in Eq.(7).
Instead of leveraging prompt tuning, ZooPFL (Lu et al., 2023) adapts the knowledge of FM to downstream clients’ tasks by training a client-specific embedding and a semantic re-mapping module that each client preserves. More specifically, each client appends the client-specific embedding to the embedding generated by a pre-trained auto-encoder, which is used to transform the client’s local input consistently with the FM. In addition, clients fine-tune encoders of their auto-encoders through FedAVG (McMahan et al., 2017) to share knowledge among them. The semantic re-mapping module aims to align the logits output from the FM with the semantic space of clients’ local tasks. ZooPFL adheres to the objective formulated in Eq. (2) with $K>1$, $\eta=0$. The domain model $\mathcal{W}^{Ck}$ of each client $k$ in ZooPFL is the client-specific embedding and semantic re-mapping module.
Representation-level knowledge transfer (RLKT) refers to transferring and adapting knowledge of FMs to domain-specific models through passing representation-level knowledge, such as intermediate forward outputs and backward gradients, between FL participating parties. In literature, split learning (Liu et al., 2022) and representation transfer are two main research directions for RLKT in federated learning.
Figure 6. Adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs) through representation-level knowledge transfer.
In a two-party split learning setting involving a client and a server, the client typically owns a private dataset and a small portion of an FM, while the server owns the larger portion. The objective of FM-based split learning is that the client collaborates with the server to train the FM using its local data, aiming to adapt the FM to its proprietary domain. More specifically, The FM is often split into an FM backbone and one or two other smaller parts. The FM backbone is the most computationally expensive part and is typically deployed in a server that has sufficient computing resources, while other parts can be deployed in either the client or server. The procedure of training the distributed FM follows the conventional deep learning procedure using stochastic gradient descent, but each portion of FM is trained by a different party (i.e., client and server).
[…]
FedBERT (Tian et al., 2022) extends the one-client setting of RAPT to the setting involving multiple clients. Clients in FedBERT first pretrain a global FM (i.e., ${\text{roberta}}_{\text{base}}$ and GPT2 (Radford et al., 2019) ) using FedAvg, and then, they collaborate with the server to train the FM backbone while each client trains its FM bottom and head locally following split learning protocol. Figure 6(a) illustrates the federated split learning with multiple clients. FedBERT corresponds to the objective formulation Eq.(2) with $K>1$, and $\eta=0$. The domain model $\mathcal{W}^{C_{k}}$ of each client $k$ is composed of an FM head and an FM bottom, and the specific loss of federated split learning $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ between the server and a client $k$ is formulated in Eq.(8).
Federated split learning enables clients to adapt the knowledge of an FM to their domain-specific tasks while consuming only a small amount of computing resources (i.e., a client only needs to train a small portion of an FM). However, clients have to collaborate with the server to make inferences, which may not be applicable to scenarios where efficient online inference is critical.
5.3. FM adapted to DMs through model-level knowledge transfer
Transferring knowledge from FMs in the form of models to downstream clients is an active research direction for grounding FMs. Model transferring methods can be reduced to the formulation of Eq.(1). We group model transferring methods into two categories. The first category is homo-model transfer, which first minimizes loss $\ell^{S}(\mathcal{W}^{S}_{\text{FM}})$ by fine-tuning or pre-training the FM $\mathcal{W}^{S}_{\text{FM}}$ and then transfers the resulting model $\mathcal{\widetilde{W}}^{S}_{\text{FM}}$ to downstream clients. $\mathcal{\widetilde{W}}^{S}_{\text{FM}}$ has the same architecture as the original FM $\mathcal{W}^{S}_{\text{FM}}$ (illustrated in Figure 7(a)). The second category is hetero-model transfer, which first minimizes loss $\ell^{S}(\mathcal{W}^{S}_{\text{FM}})$ by compressing the FM $\mathcal{W}^{S}_{\text{FM}}$ and then transfer the resulting model $\mathcal{\widetilde{W}}^{S}_{\text{SM}}$ to downstream clients. $\mathcal{\widetilde{W}}^{S}_{\text{SM}}$ has a different architecture as the original FM $\mathcal{W}^{S}_{\text{FM}}$ (illustrated in Figure 7(b)). A model transferring method involves domain adaptation if $\lambda>0$ and privacy protection if $\eta>0$. In the following, we review the relevant literature.
@@ 5.3. FM adapted to DMs through model-level knowledge transfer
Figure 7. Transferring knowledge of the server’s foundation model (FM) to clients through model transfer.
@@
For transferring hetero-models adaptive to downstream tasks,  Wang et al. (2023) proposed an approach that leverages public data and pre-trained FM (i.e., LaMDA (Cohen et al., 2022)) to enhance downstream cross-device federated learning. More specifically, they first leverage their proposed distribution matching algorithm to sample public data closely resembling the distribution of FL clients’ private data, and then they use the sampled data to distill a pre-trained FM (i.e., LaMDA) into an on-device language model, which serves as an initialization for FL clients. GPT-FL (Zhang et al., 2023a) leverages FMs (i.e., Stable Diffusion (Rombach et al., 2022), SpeechT5 (Ao et al., 2022), and AudioLDM (Liu et al., 2023a)) to generate diversified synthetic data to train a downstream model, which is then distributed to FL clients for model initialization.

### 29 Dec 2023 v7

@@ 3.2. Formulations of Machine Learning Tasks
where $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is the local FM of client $k$, $\phi_{\mathcal{W}^{G}_{a}}$ is the global PEFT module (e.g., prompt model, adaptor, and LoRA) parameterized by $\mathcal{W}^{G}_{a}$, and $\oplus$ denotes an operator that composes $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ with PEFT module $\phi_{\mathcal{W}^{G}_{a}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for optimizing the global PEFT module $\phi_{\mathcal{W}^{G}_{a}}$ based on client $k$’s local data $x^{C_{k}}$ and $y^{C_{k}}$ while keeping $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ frozen.
The research works, FedPETuning (Zhang et al., 2023d), FedPrompt (Zhao et al., 2023b), FedAdapter (Cai et al., 2023b), and FedIT (Zhang et al., 2023c) adhere to the formulation of Eq.(13). FedPrompt, FedAdapter, and FedIT specifically investigate the impact of prompt tuning, adapter, and LoRA under federated learning, respectively. On the other hand, FedPETuning delves into exploring four PEFT techniques, namely adapter tuning, prefix tuning, LoRA, and BitFit, while operating under the FedAvg protocol.
@@ 4. FTL-FM Taxonomy
Semi-Honest $\mathcal{D}^{S}$ ($\mathcal{W}^{S}$) PPDG  (Carranza et al., 2023; Kurakin et al., $*{}^{*}$  (Mattern et al., 2022; Yue et al., $*{}^{*}$ DP  (Behnia et al., 2022; Bu et al., 2022; Shi et al., $*{}^{*}$, MP (Youssef et al., $*{}^{*}$ DP  (Yu et al., 2022; Mireshghallah et al., $*{}^{*}$ CP (Xiao et al., 2023)  (Fan et al., 2023; Chua et al., 2023; Khalid et al., ${}^{\dagger}$
$\mathcal{D}^{C}$ ($\mathcal{W}^{C}$) DP (Li et al., 2023c; Shen et al., $*{}^{*}$, WPE (Xu et al., $*{}^{*}$, PTB (Zhou et al., $*{}^{*}$ PPDS (Hou et al., 2023b; Yu et al., $*{}^{*}$ DP (Wang et al., $*{}^{*}$ SA (Fan et al., $*{}^{*}$ DP (Azam et al., 2023; Balunovic et al., $*{}^{*}$, FE (Gupta et al., $*{}^{*}$ HE (Jin et al., $*{}^{*}$ DP  (Xu et al., 2023a; Zhao et al., $*{}^{*}$, PEFT (Zhang et al., $*{}^{*}$, SA (Fan et al., $*{}^{*}$
@@
5. FTL-FM: Foundation Models for Enhancing Domain Models
In this section, we overview FTL-FM methods that fall into setting 1⃝ (see Definition 3.1), the objective of which is to transfer and adapt the knowledge of foundation models (FMs) to downstream clients’ domain models (DMs) for exploiting the power of FMs. To this end, the literature proposed a variety of methods involving data-level knowledge transfer, representation-level knowledge transfer, and model-level knowledge transfer.
5.1. FM adapted to DMs through data-level knowledge transfer
The literature has explored mainly two categories of methods that adapt the knowledge of FMs to specific domains through data-level knowledge transfer: (1) An FM server transfers data generated by general-purpose FMs or FMs augmented with industry-level knowledge (e.g., LawGPT (Nguyen, 2023), FinGPT (Yang et al., 2023), BloombergGPT (Wu et al., 2023a), and Med-PaLM2 (Singhal et al., 2023)) to clients of specific domains; (2) a downstream client prompts tuning a closed-source FM (e.g., ChatGPT) by learning a local prompt generator based on domain-specific data and predictions generated by the FM. We summarize data-level knowledge transfer methods that fall into setting 1⃝ in Table 3222While we consider an FTL-FM approach should incorporate certain protection mechanisms to protect data or models, we still review highly relevant works that prioritize the development of knowledge transfer methods without considering data or model protection and include them in the Table to ensure thoroughness. (same for Section 6 and Section 7). These methods leave ”blank” in the ”What To Protect” and ”How To Protect” columns. and elaborate on these methods in this subsection.
@@ 5.1. FM adapted to DMs through data-level knowledge transfer
BBT (Sun et al., 2022b, a) $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + soft $P$
@@
BDPL (Diao et al., 2023) $\hat{Y}$ from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + hard $P$
DSP (Li et al., 2023b) Reward from $\mathcal{W}^{S}_{\text{FM}}$ $X$ + hard $P$
Fed-BBPT (Lin et al., 2023)
Fr. $S_{\text{FM}}$:
$\hat{Y}$ of $\mathcal{W}^{S}_{\text{FM}}$
Fr. $S_{\text{Fed}}$:
Global model
To $S_{\text{FM}}$:
$X$ + soft $P$
To $S_{\text{Fed}}$:
$\mathcal{W}^{C}$
FedBPT (Sun et al., 2023)
Fr. $S_{\text{FM}}$:
$\hat{Y}$ of $\mathcal{W}^{S}_{\text{FM}}$
Fr. $S_{\text{Fed}}$:
Global distr. of $P$
To $S_{\text{FM}}$:
$X$ + soft $P$
To $S_{\text{Fed}}$:
Local distr. of $P$
ZooPFL (Lu et al., 2023)
Fr. $S_{\text{FM}}$:
$\hat{Y}$ of $\mathcal{W}^{S}_{\text{FM}}$
Fr. $S_{\text{Fed}}$:
Global encoder
To $S_{\text{FM}}$:
Perturbed $X$
To $S_{\text{Fed}}$:
Local encoder
Figure 6. Adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs) through data-level knowledge transfer. (a) illustrates that the knowledge of FM is transferred to downstream clients via privacy-preserving synthetic data sharing. (b) illustrates that a client can adapt knowledge of the FM to its local task via federated prompt learning.
Alternatively, clients can adapt knowledge of the FM hosted by an FM server to their specific domains through federated prompt tuning. A typical setting of federated prompt tuning involves a client that learns a local model to generate prompts that can guide the FM toward the desired responses (illustrated in Figure 6(b)). For example, BBT (Sun et al., 2022b), BBTv2 (Sun et al., 2022a), and BlackVIP (Oh et al., 2023) proposed that a client can customize a remote FM (e.g., ${\text{RoBERTa}}_{\text{large}}$ (Liu et al., 2019), ${\text{BART}}_{\text{large}}$ (Lewis et al., 2020), and ViT-B (Radford et al., 2021)) by optimizing a domain-specific prompt generator based on its local data together with the predictions (e.g., logits) generated by the remote FM. BBT and BBTv2 employ a gradient-free method called CMA-ES (Covariance Matrix Adaptation Evolution Strategy) (Hansen, 2016) to optimize the client’s local prompt generator, whereas BlackVIP leverages SPSA (Simultaneous Perturbation Stochastic Approximation) (Spall, 1992) to efficiently approximates the high-dimensional gradients, which then are used to optimize the prompt generator. BBT, BBTv2, and BlackVIP require the FM API to take continuous (i.e., soft) prompts as input, and thus, they cannot be applied to FM APIs that only accept discrete inputs. To eliminate this constraint, BDPL (Diao et al., 2023) and DSP (Li et al., 2023b) were proposed to learn a client side’s discrete (i.e., hard) prompt generator (DPG) for customizing FM’s responses. BDPL adopts a variance-reduced policy gradient algorithm (Zhou et al., 2021) to optimize the DPG, while DSP optimizes the DPG through supervised fine-tuning and reinforcement learning. BBT, BBTv2, BlackVIP, BDPL, and DSP involve only one client, and thus, they correspond to stage 2 of the first objective formulated in Eq.(2) with $K=1,p^{C_{1}}=1$, $\eta=0$, the domain model $\mathcal{W}^{C_{1}}$ is the client’s local prompt generator and the specific form of the loss $\ell^{C_{1}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{1}};d)$ is formulated in Eq.(7).
Fed-BBPT (Lin et al., 2023) and FedBPT (Sun et al., 2023) extend the federated prompt tuning with only one client to multiple clients, the workflow of which is illustrated in Figure  7. More specifically, Fed-BBPT and FedBPT involve two servers: an FM server, denoted as $S_{\text{FM}}$, and a federated server, denoted as $S_{\text{Fed}}$. The former hosts an FM to be prompted by clients, while the latter is for federated tuning clients’ local prompt generators. In Fed-BBPT, each client leverages the Simultaneous Perturbation Stochastic Approximation (Spall, 1997) to train its prompt generators based on local data and predictions generated by the FM, and the federated server aggregates clients’ local prompt generators to form the global prompt generator. FedBPT differs from Fed-BBPT in that each client in FedBPT exploits CMA-ES to optimize the distribution of prompts generated by its prompt generator, and the server derives the global prompt generator by aggregating local distributions of clients’ prompts. Generally, Fed-BBPT and FedBPT adhere to the objective formulated in Eq.(2) with $K>1$, $\eta=0$, the domain model $\mathcal{W}^{C_{k}}$ of each client $k$ is a prompt generator and the specific loss of federated prompt tuning $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ between the FM server and a client $k$ is formulated in Eq.(7).
Figure 7. Adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs) through data-level knowledge transfer. This figure illustrates that multiple clients adapt the knowledge of the FM hosted by an FM server to their local tasks via federated prompt tuning, and these clients also leverage the FedAvg algorithm or its variants with the help of the federated server to train their local domain models.
Instead of leveraging prompt tuning, ZooPFL (Lu et al., 2023) adapts the knowledge of FM to downstream clients’ tasks by training a client-specific embedding and a semantic re-mapping module for each client. More specifically, each client appends the client-specific embedding to the embedding generated by a pre-trained auto-encoder, which aims to make the client’s local input consistent with the FM. Clients also collaboratively fine-tune encoders of their auto-encoders through FedAVG (McMahan et al., 2017) to share knowledge. The semantic re-mapping module aligns the logits output from the FM with the semantic space of clients’ local tasks. ZooPFL adheres to the objective formulated in Eq. (2) with $K>1$, $\eta=0$. The domain model $\mathcal{W}^{Ck}$ of each client $k$ in ZooPFL is the client-specific embedding and semantic re-mapping module.
Representation-level knowledge transfer (RLKT) refers to transferring and adapting knowledge of FMs to domain-specific models through passing representation-level knowledge, such as intermediate forward outputs and backward gradients, between FL participating parties. In literature, split learning (Liu et al., 2022) and representation transfer are two main categories of RLKT explored in the setting 1⃝ of FTL-FM. We summarize the two categories of RLKT methods in Table 4 and elaborate on these methods in this subsection.
[…]
$\mathcal{D}^{S}$ $\mathcal{W}^{S}_{\text{FM}}$ $\mathcal{D}^{C}$ $\mathcal{W}^{C}$
@@ 5.2. FM adapted to DMs through representation-level knowledge transfer
[…]
FedBERT (Tian et al., 2022) Forward Repr, Backward gradient Forward Repr, Backward gradient
@@
In a two-party federated split learning setting involving a client and a server, the client typically owns a private dataset and a small portion of an FM, while the server owns the larger portion. The objective of FM-based split learning is to adapt the knowledge of FM to the client’s specific domain by fine-tuning the distributed FM using the client’s local data. The FM is often split into an FM backbone and one or two other smaller parts. The FM backbone is the most computationally expensive part and is typically deployed in a server with sufficient computing resources, while other parts can be deployed in either the client or server. The training procedure of distributed FM follows the conventional split learning procedure using stochastic gradient descent, but each portion of FM is trained by a different party (i.e., client and server).
Figure 8. Adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs) through representation-level knowledge transfer. (a) illustrates that the knowledge of FM is adapted to clients’ DMs through federated split learning, while (b) through federated representation transfer.
[…]
FedBERT (Tian et al., 2022) extends the federated split learning with one client to multiple clients, the workflow of which is illustrated in Figure 8(a). In FedBERT, a pre-specified FM (e.g., ${\text{roberta}}_{\text{base}}$ and GPT2 (Radford et al., 2019)) is split into three parts: FM backbone, FM head, and FM bottom. The FM backbone is deployed on the server, while the FM head and FM bottom are deployed on each client. Clients in FedBERT collaborate with the server to train the FM backbone and their corresponding FM bottoms and heads following conventional split learning protocol (Liu et al., 2022). FedBERT adheres to the objective formulation Eq.(2) with $K>1$, and $\eta=0$. The domain model $\mathcal{W}^{C_{k}}$ of each client $k$ is composed of an FM head and an FM bottom, and the specific loss of federated split learning $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ between the server and a client $k$ is formulated in Eq.(8).
Federated split learning enables clients to adapt the knowledge of an FM to their domain-specific tasks while consuming only a small amount of computing resources (i.e., a client only needs to train a small portion of an FM). However, clients have to collaborate with the server to make inferences, which may not apply to scenarios where efficient online inference is critical.
5.3. FM adapted to DMs through model-level knowledge transfer
Transferring knowledge from FMs in the form of models to downstream clients is an active research direction for grounding FMs. Model transferring methods can be reduced to the formulation of Eq.(1). We group model transferring methods into two categories. The first category is homo-model transfer, which first minimizes loss $\ell^{S}(\mathcal{W}^{S}_{\text{FM}})$ by fine-tuning or pre-training the FM $\mathcal{W}^{S}_{\text{FM}}$ and then transfers the resulting model $\mathcal{\widetilde{W}}^{S}_{\text{FM}}$ to downstream clients. $\mathcal{\widetilde{W}}^{S}_{\text{FM}}$ has the same architecture as the original FM $\mathcal{W}^{S}_{\text{FM}}$ (illustrated in Figure 9(a)). The second category is hetero-model transfer, which first minimizes loss $\ell^{S}(\mathcal{W}^{S}_{\text{FM}})$ by compressing the FM $\mathcal{W}^{S}_{\text{FM}}$ and then transfer the resulting model $\mathcal{\widetilde{W}}^{S}_{\text{SM}}$ to downstream clients. $\mathcal{\widetilde{W}}^{S}_{\text{SM}}$ has a different architecture as the original FM $\mathcal{W}^{S}_{\text{FM}}$ (illustrated in Figure 9(b)). A model transferring method involves domain adaptation if $\lambda>0$ and privacy protection if $\eta>0$. We summarize the two categories of model-level knowledge transfer methods in Table 5 and illustrate them in Figure 9(a) and Figure 9(b), respectively. In the following, we will elaborate on these methods of the two categories.
[…]
$\mathcal{D}^{S}$ $\mathcal{W}^{S}_{\text{FM}}$ $\mathcal{D}^{C}$ $\mathcal{W}^{C}$
@@ 5.3. FM adapted to DMs through model-level knowledge transfer
GPT-FL (Zhang et al., 2023b) Model trained on data generated by $\mathcal{W}^{S}_{\text{FM}}$
@@
BD-FMFL (Li et al., 2023e) BD global model
@@
Figure 9. Transferring knowledge of the server’s foundation model (FM) to clients through model transfer. (a) illustrates the knowledge of FM is transferred to downstream clients through homo-model transfer, while (b) through hetero-model transfer.
@@
For transferring hetero-models adaptive to downstream tasks,  Wang et al. (2023) proposed an approach that leverages public data and pre-trained FM (i.e., LaMDA (Cohen et al., 2022)) to enhance downstream cross-device federated learning. More specifically, they first leverage their proposed distribution matching algorithm to sample public data closely resembling the distribution of FL clients’ private data, and then they use the sampled data to distill a pre-trained FM (i.e., LaMDA) into an on-device language model, which serves as an initialization for FL clients. GPT-FL (Zhang et al., 2023b) leverages FMs (i.e., Stable Diffusion (Rombach et al., 2022), SpeechT5 (Ao et al., 2022), and AudioLDM (Liu et al., 2023a)) to generate diversified synthetic data to train a downstream model, which is then distributed to FL clients for model initialization.